# Idea-Gated Transformers: Enforcing Semantic Coherence via Differentiable Vocabulary Pruning

### v1

@@ 7 Conclusion
We presented the Idea-Gated Transformer, a novel architecture that enforces semantic coherence via differentiable vocabulary pruning. By training an auxiliary head to predict future semantic content and using it to gate the primary generation head, we successfully decoupled “planning” from “speaking.”
Our experiments on WikiText-103 demonstrate that this mechanism significantly reduces topic drift. In specialized domains such as Chemistry and Hardware, the Gated model exhibited up to 50% higher domain retention compared to a standard GPT baseline (radford2019language). Qualitative analysis confirms that the gating mechanism actively suppresses out-of-domain associations, forcing the model to maintain narrative focus even over long generation windows.
While the current implementation requires careful tuning of the gating strength ($\alpha$) to avoid repetition loops, it offers a promising proof-of-concept for Controllable Generation. As LLMs are increasingly deployed in high-stakes domains requiring strict adherence to topic and logic, architectural constraints like Idea Gating provide a robust alternative to purely prompt-based engineering (ouyang2022training).

### v2

@@ 7 Conclusion
We presented the Idea-Gated Transformer, a novel architecture that enforces semantic coherence via differentiable vocabulary pruning. By augmenting a frozen Mistral-7B backbone with a lightweight auxiliary planning head, we successfully decoupled the cognitive processes of “planning” (System 2) and “speaking” (System 1).
Our experiments on FineWeb-Edu demonstrate that this mechanism solves the fundamental problem of Associative Drift without degrading modeling performance. The Idea-Gated model achieved a superior validation perplexity compared to a standard QLoRA baseline (7.78 vs. 8.08) while exhibiting superior qualitative control (dettmers2023qlora). In adversarial stress tests, the gating mechanism successfully suppressed high-probability hallucinations (e.g., stopping the “Batman” drift) and resolved complex polysemy where the baseline failed (holtzman2019curious).
This work suggests that the limitations of Next-Token Prediction are not inherent to the Transformer architecture but to the objective function itself (brown2020language). By explicitly modeling future ideas as a constraint on present words, we offer a robust, parameter-efficient path toward Large Language Models that think before they speak.